a coin is tossed 4 times and the sequence of heads and tails is observed:
(a)how many different outcomes are possible
(b)how many different outcomes have exactly 3 tails
(c)how many different outcomes have at least 1 tail
Please Please help me with this-thank you very very much

Answer by stanbon(57361)

a coin is tossed 4 times and the sequence of heads and tails is observed:
(a)how many different outcomes are possible
Ans: 2^4 = 16 outcomes
---------------------------------
(b)how many different outcomes have exactly 3 tails
Ans: Same as "how many have exactl one head": that number is 4
--------------------------------------------------------------------
(c)how many different outcomes have at least 1 tail
#(at least 1 tail) = 16 - #(no tail) = 16-1 = 15
======================================================
Cheers,
Stan H.